OpenStudy (anonymous):

if f(x)=2x+8 and g(x)=x^4, what is (g of f)(-3)?

OpenStudy (shayaan_mustafa):

Hi JGAZAY7 :) How are you?

OpenStudy (shayaan_mustafa):

It is not degree my friend. It is read as, g of f. OK

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

@shayaan_mustafa

OpenStudy (shayaan_mustafa):

g of f means, g(f(x)) This means. Put the function of f inside the g. like this, \(g(f(x))=(2x+8)^4\)

OpenStudy (shayaan_mustafa):

Now put -3 for this composite function. like this, \[g(f(-3)) = (2(-3)+8)^4\] \[g(f(-3)) = (-6+8)^4\] \[g(f(-3)) = (2)^4\] \[g(f(-3)) = 16\]

OpenStudy (anonymous):

Thanksssssss

OpenStudy (shayaan_mustafa):

Welcome buddy.